OpenStudy (anonymous):

If the velocity is 2.794×103 m/s and the period is 2.083×10-1 s, what is the wavelength?

OpenStudy (anonymous):

T=1/f v=f*lmda

OpenStudy (anonymous):

wavelength= v/(1/T)

OpenStudy (anonymous):

Yes, v = f lambda, where f is frequency and lambda is the wavelength 3794 m/s = (1/.2083) lambda lambda = 3794 x .2083 = 790 m

OpenStudy (anonymous):

Time period(T)=1/f 2.083×10-1 = 1/f Frequency=1/2.083×10-1 Wavelength=v/f=(2.791×103/1/2.083×10-1) =2.791×103×2.083×10-1 =581.3653m